If scores are normally distributed with a mean of 30 and a standard deviation of 5.what percentage of the scores is?

1
Expert's answer
2022-04-27T14:21:47-0400

If scores are normally distributed with a mean of 30 and a standard deviation of 5, what percent of the scores is:

(a) greater than 30?

(b) greater than 37?

 (c) between 28 and 34?

(a)


"P(X>30)=1-P(Z\le\dfrac{30-30}{5})=0.5"


50%


(b)


"P(X>37)=1-P(Z\le\dfrac{37-30}{5})"

"=1-P(Z\le1.4)\approx0.080757"

8.0757%



(c)


"P(28<X<34)=P(Z<\dfrac{34-30}{5})"

"-P(Z\le\dfrac{28-30}{5})=P(Z<0.8)-P(Z\le-0.4)"

"\approx0.7881446-0.3445783\approx0.443566"

44.3566%